A 3500-kg car hits an 80-kg pedestrian who is standing in the middle of the road. Before the collision, the car was travelling at 30 km/h and the pedestrian was stationary. After the collision, the car was traveling at 28.5 km/h. At what speed will the pedestrian be flung down the road (in km/h)


1
Expert's answer
2022-01-02T18:29:04-0500

mass of Car "(M_c)=3500kg"

Mass of pedestrian"(M_p)=80kg"

Initial velocity before collision


"V_c=30km/hr=\frac{30\times1000}{3600}m/sec=\frac{150}{18}m/sec\\v_p=0km/hr"

After collision


"V'_c=28.5km/hr=\frac{28.5\times1000}{3600}m/sec=\frac{1425}{180}m/sec"

Momentum conservation

Momentum Before collision= momentum after collision

"M_cV_c+M_pV_p=M_cV'_c+M_pV'_p"

Put value


"3500\times\frac{150}{18}+80\times0=3500\times\frac{1425}{180}+80\times V'_p"

"29166.66=27708.33+80V'_p"

"1458.33=80V'_p\\V'_p=28.2291m/sec"